Good Luck. (Need to show all your work to get full credit; no work, no credit) 1.For the given sample data below: a Construct a frequency distribution using six classes for a sample data that includes. Class Limits (5 pts), Class Boundaries (6 pts), Frequency (3 pts), Relative frequency (3 pts), Cumulative Frequency (3 pts), Midpoint X.5 pts), f-x-(5 pts), f. (6 pts). b. For constructed grouped data find: mean (8 pts), mode for grouped data (2 pts), modal class (2 pts), the variance (10 pts), and the standard deviation (2 pts) c. Draw histogram and discuss the shape of a distribution (10 pts), frequency polygon (10 pts) ogive (10 pts), and pie chart for the following table (10 pt). NOTE: you MUST use 2.0 as a lower limit of the first class and do appropriate estimation for class width and in other computations. 4.6 3.9 2.8 6.6 4.2 3.7 3.7 5.9 3.2 2.2 3.2 4.1 21 4.2 50 46 54 24 63 29 5.3 40 4.7 3.6 3.3 6.9 45 2.5 5.4 5.7 3.84.1 7.96.2 3.0 33 5.0 5.4 3.4 4.4 4.0 365.0 4.1 48 72 64 30 3.5 53 7.7 3.9 26 5.6 33 55 43 3.9 63 3.1 3.1 3.0 4.8 4.1 5 Class Limits Class boundaries Tally Frequency Relative CumuativeClas Frequency regueny
Coneegnydbutionsing six classes for a sample data that includes. Cess umits i pi Cls Bndries i6 pts,Frequency (3 pts), Relative frequency (3 pts) ๛ for co-ted Enen dat. find mean (8 pts), mode for grouped data (2 pts), modal class (2 pts), the (10 pts), and the standard deviation (2 pts) Orew histogram and dissvss the shape of a distribution (10 pts), frequency polygon (10 pts). give (30 pts), and pie chart for the following table (10 pt). NOTE: you MUST use 2.0 as a lower limit of the first class and do appropriate estimation for class width and in other computations. 2.4 3.0 3.3 0 4.1 Tally Feeguency Frequency idpoin You must show ALL your computations.
Answer #1

a.

Class limitClass boundaries Frequency (f) Relative frequency Cumulative frequencvl Mid value(xm) f*xm 2.0-2.9 3.0-3.9 4.0-4.9 5.0-5.9 6.0-6.9 7.0-7.9 f*xmA2 1.95-2.95 2.95-3.95 3.95-4.95 4.95-5.95 5.95-6.95 6.95-7.95 Total 17.1542.0175 3.45 6555 226.1475 71.2 316.84 65.41 356.43 38.7 249.615 22.35 166.5075 280.35 1357.558 2.45 19 16 26 42 0.3016 0.254 0.1905 0.0952 0.0476 4.45 5.45 6.45 63 63

b. Mean=280.35/63=4.45

Modal class=2.95-3.95

Mode+ m where, xi lower class boundary of modal class 2.95 fm-frequency of modal class 19 fm-1 = frequency of preceding class

(c)

7.9 2.1 5.8 Maximum Minimum Range No. of classes- Class width- Class limit Class boundaries Frequcy(f) Relative frequency CumHence shape of the distribution is positively skewed.